OpenStudy (anonymous):

25mL of ethanol (C2H3OH) is diluted to a volume of 150mL with water. What is the concentration (in % v/v) of the following solution?

OpenStudy (zepp):

Total: 25+150 = 175ml Percentage V/V = 25/175 * 100 \(\approx\) 14.2857%

OpenStudy (zepp):

It mean that 14.2857% of every 175ml of this solution is ethanol.

OpenStudy (zepp):

Like beers, e.g.: 5% alcool mean that there's 5ml of pure alcohol in 100ml of water. So there's 5ml of alcohol and 95ml of water.

OpenStudy (anonymous):

That clarifies a lot. Thanks so much!

OpenStudy (zepp):

You are welcome :)

OpenStudy (zepp):

Forget my blahblah, I got mixed up again.

OpenStudy (zepp):

Concentration % V/V = Quantity of the product / Quantity of the solution

OpenStudy (zepp):

There Quantity of the product : 25ml Quantity of the solution: 150ml + 25ml = 175 Sorry for the confusion :(

OpenStudy (anonymous):

No big deal! thank you for the help!